# Does rare, noise-induced, bypass transition in plane Couette flow bypass instantons ?

### v1

@@ 1 Introduction
Many physical systems, after they underwent a subcritical bifurcation or a first order transition (or a similar transition in more complex systems), display two or more metastable states. This situation means that the physical system can switch from one metastable state to another if it undergoes some statistically steady forcing, usually by crossing a boundary separating the basin of attraction of one state from another. Conversely, this physical system will relax toward either states if it is prepared with an initial condition on the corresponding side of the separatrix. Traditional bistable systems are found in kinetic chemistry [Onsager, 1938, Hänggi et al., 1990, Van Erp et al., 2003, Lopes and Lelièvre, 2019], where the two states are the reactants and the products or different configurations or conformations [Lopes and Lelièvre, 2019] of a complex molecule, while the force driving transitions is the thermal noise. Much insight on bistability has been gained from formal noise driven systems [Metzner et al., 2006, Cérou et al., 2011, Bouchet and Touchette, 2012, Rolland and Simonnet, 2015, Rolland et al., 2016, Lucente et al., 2022]. More recently, a lot of effort has also been directed toward the study of bistable climate systems [Lucarini and Bódai, 2017, Lucarini and Bódai, 2019, Baars et al., 2021, Herbert et al., 2020] and bistable turbulent flows [Podvin and Sergent, 2017] of aerodynamical [Grandemange et al., 2013, Kim and Durbin, 1988] or geophysical interest [Berhanu et al., 2007, Bouchet et al., 2019]. In these systems, the metastable states are usually large scale coherent circulations, like jets, or vortices, while the force driving transition is intrinsic to the system and related to the fluctuating small scales of turbulence.
@@
A central object of study for metastable systems is the succession of states visited by the system when it transits from one bistable state to another. In kinetic chemistry, this part of the dynamics is termed a reactive trajectory [Hänggi et al., 1990] or a transition path [Van Erp et al., 2003]. In kinetic chemistry, and more generally, in a wide class of stochastic systems, as the thermal noise is decreased and the transition becomes rare, the transition paths tend to concentrate around a specific path termed the instanton. The instanton generally consists in a fluctuation path under noise toward a specific transition state which is a saddle of the deterministic part of the dynamics, followed by a deterministic relaxation path [Touchette, 2009] (Fig. 15 (a)). This path is important because it highlights the physical mechanisms of the transition, while the properties of the transition state control the parametric dependence of the transition rate [Hänggi et al., 1990, Touchette, 2009, Bouchet and Reygner, 2016]. While the instanton always visits the transition state, complex trajectories can be found in non-gradient systems [Wan et al., 2015]. The rarity and the concentration of trajectories is quantitatively studied in the case where noise variance is a constant times $\epsilon$, a vanishing parameter. One then finds that the transition rate, or its inverse the mean first passage time before a transition occurs $T$, varies exponentially with $\epsilon$, as
@@
where $\mathcal{I}$ is termed a rate function: it is finite (the limits are defined), not everywhere zero and is independent on $\epsilon$. This rate function is directly controlled by the properties of the escaped multistable state and the saddle. In kinetic chemistry, for relatively low temperatures, this relation was identified as Arrhenius law. It was then formalised in the context of stochastic gradient systems as Eyring–Kramers theory [Hänggi et al., 1990]. Such a formula can also be obtained in non gradient systems [Bouchet and Reygner, 2016]. In general, the asymptotic study of exponential dependences in probabilities or rate of probabilities (Eq. (1)) is termed Large Deviations, and encompasses the case of rare reactive trajectories. In the case of rare paths, this Large Deviation principle is named after Freidlin and Wentzell [Freidlin and Wentzell, 1998]. The difficulty of computing the transition path directly from the dynamics, either because of its complexity or because direct simulation times are too long for molecular dynamics, has led to the development and application of rare event simulation methods for kinetic chemistry [Van Erp et al., 2003, Lopes and Lelièvre, 2019]. Another approach for these computations consists in exploiting the fact that the instanton minimises the so-called Onsager–Machlup action [Onsager and Machlup, 1953] in optimisation procedures [Grafke and Vanden-Eijnden, 2019, Wan, 2013, Wan et al., 2015, Wan and Yu, 2017, Börner et al., 2023]. Note that the noise covariance matrix is involved in the action, so that the type of forcing can influence the fluctuation path and the transition rate. The transition state, however, is a property of the deterministic part of the dynamics and thus has a degree of universality.
@@
The success of rare event simulation methods in chemistry has encouraged their use in fluid dynamics. For that purpose, they have been applied to spatially extended systems of gradual complexity, starting from one dimensional stochastically forced systems [Rolland, 2018, Baars et al., 2021], moving to two dimensional stochastically forced turbulence [Bouchet et al., 2019] and recently reaching three dimensional deterministic transitional turbulence [Rolland, 2022, Gomé et al., 2022]. The progress has been milestoned in this manner because of the size of the leap between stochastic differential equation and fully developed turbulent flows. This means a drastic increase of the number of degrees of freedom (DoF) involved, a complete change of the force driving the metastability, and sometimes a strong increase in complexity of the transition state between the two metastable states. While the extended systems studied were firstly chosen for their physical interest, the order of studies was also motivated by the need to tackle this changes separately from one another at first.
@@
This was the case of the most recent works, concerning the collapse of turbulence in plane Couette flow [Rolland, 2022, Gomé et al., 2022]. Plane Couette flow (PCF) is the flow between two parallel moving walls (Fig. 1, left). In PCF, like in other wall flows such as plane Poiseuille flow [Wan, 2013, Wan et al., 2015, Wan and Yu, 2017], Couette–Poiseuille flow [Liu et al., 2021], Hagen–Poiseuille pipe flow [Willis and Kerswell, 2009] or boundary layers [Spangler and Wells Jr, 1968, Rigas et al., 2021], wall turbulence can coexist at Reynolds number (or positions in the case of boundary layers) for which the laminar baseflow is linearly stable (see [Romanov, 1973] for the case of PCF). As a consequence, transitional turbulence can collapse under its own fluctuations, while the flow can go from laminar to turbulent under a forcing [Spangler and Wells Jr, 1968, Rigas et al., 2021] or if it is given an initial condition “close enough” to turbulence [Schmiegel and Eckhardt, 1997]. The study of the collapse of turbulence using rare event methods gave the occasion to investigate the physical mechanism of laminar hole opening as well as the technical solutions to study rare multistability events in a three dimensional flow not stochastically forced [Rolland, 2022, Gomé et al., 2022]. Conversely, the study of the forcing of laminar flow toward turbulence, under a noise or another, would give the occasion to investigate the crossing of a very complex separatrix [^1] We use the term separatrix for lack of a better word. strictly speaking, an unforced wall flow starting from any initial condition eventually reaches the laminar state: what changes is the relaxation time before it does so. This duration dramatically increases as an hypersurface is crossed [Schmiegel and Eckhardt, 1997]..
@@
As such, this has brought the question of the separatrix between laminar and turbulent flow at the forefront of the study of transition. The first descriptions by Waleffe and collaborators of the mechanisms maintaining transitional wall turbulence was the self sustaining process (SSP) of wall turbulence [Hamilton et al., 1995, Waleffe, 1997]. This process represented wall turbulence as a stable limit cycle, which was therefore separated from the laminar baseflow in phase space by an actual separatrix with a saddle on it. The SSP describes how streamwise velocity streaks and streamwise vortices regenerate one another. It can take place provided the Reynolds number is large enough and the domain size is larger than the so called Minimal Flow Unit (MFU) [Jiménez and Moin, 1991]. Refinements of that picture indicated that this process was not periodic nor an attractor (thus leaving the room for collapse of turbulence) but nevertheless confirmed the idea that there was a separatrix, on which saddles could be found. The structure of the separatrix, which is a key matter for the forced path from laminar to turbulent flow, was often studied using methods of dichotomy or bisections [Nusse and Yorke, 1989]. This has led to the study of edge states, saddles on that separatrix, in wall flows [Toh and Itano, 2003, Schneider et al., 2007, Schneider et al., 2008, Schneider et al., 2010, Willis and Kerswell, 2009] as well as in climate models [Lucarini and Bódai, 2017]. Note that additional unstable limit cycles and fixed points can be found on the turbulent side of the separatrix [Gibson et al., 2008]. In wall flows, these unstable fixed points and limit cycles are often reminiscent of the turbulent flow, in that they display velocity streaks and streamwise vortices. However, they display less small scale fluctuations and are often very spatially regular. This regularity can be characterised using the symmetry groups of the flow [Gibson et al., 2008, Schneider et al., 2010]. Note finally that, as the separatrix and the unstable states on it are a property of the flow in the configuration of interest and not of initial conditions or forcing type, they also represent an universal feature of transition to turbulence.
@@
In wall flows, the complexity of this boundary is such that the saddle on it is not the point of the separatrix which is has the smallest amplitude of departure to the laminar baseflow. Experimental and numerical studies of initial conditions, generated using various specific rules, of minimal energy that evolve into turbulence have highlighted that. This leaves the question of whether the most probable forced path would visit this saddle at very small but finite forcing noise variance. Alternate investigations have thus been aimed at identifying a path forced and/or triggered by a finite amplitude initial condition that would drive the flow from laminar to turbulent. Linear studies of optimal initial conditions of infinitesimal amplitude leading to maximal amplification of energy identified lift-up: the extraction of energy from the laminar base flow from streamwise vortices. The lift-up mechanism was thus identified in transient growth study that highlight the fact that energy growth can happen while all linear modes are stable is the linearised Navier–Stokes equations lead to a non normal operator [Farrell and Ioannou, 1993a, Schmid and Henningson, 1994, Trefethen et al., 1993]. Such an optimal (in shape) initial condition whose amplitude places it on the laminar side of the boundary would remain on the laminar side of said boundary. However, for initial conditions starting on the turbulent side of the boundary, lift-up is an important part of the self-sustaining process [Waleffe, 1997]. Refinement of optimisation of initial condition energy amplification technics can be used to compute the optimal finite amplitude initial condition leading to maximal non-linear energy growth [Cherubini et al., 2011, Rabin et al., 2012, Monokrousos et al., 2011]. This initial condition is termed a minimal seed. Such computations are still performed on the laminar side of the separatrix, however a slight modification of the seed, in the right direction, can generate an initial condition which has very little energy but can nevertheless travel to turbulence extremely efficiently.
@@
The question of whether a random forcing of a wall flow can generate a minimal seed, a flow organisation whose shape and amplitude is precisely defined, remains open. Indeed, the study of forced wall flows gives little weight to effects like the Orr mechanism, involved in the growth of minimal seeds. In parallel with studies of optimal initial conditions leading to transient growth, the study of the response of the linearised Navier–Stokes equations to a forcing with a wide range of spatial spectra, which are white in time, identified a reacting flow structure [Farrell and Ioannou, 1993b]. The weakly non linear behaviour of flows under such stochastic forcing can also be performed [Ducimetière et al., 2022]. The inclusion of the strong non-linearities may very well strongly change this picture because they introduce the aforementioned complex separatrix. The experimental study of the response of boundary layers to harmonic forcing of different spectrum has a long history and shows a strong dependence on the forcing frequency and amplitude of the point of appearance of wall turbulence [Spangler and Wells Jr, 1968]. Numerical equivalent of such studies of the response to an harmonic forcing have only been recently undertaken [Rigas et al., 2021]. Meanwhile, to refocus on our subject of interest, studying the response to a stochastic forcing of vanishing variance requires a rare event simulation method. This has been performed in the case of a model of Hagen–Poiseuille pipe flow [Rolland, 2018]. In that case, we have termed the forcing of the flow model from laminar to turbulent build up. However the separatrix in that model is nowhere near as complex as the one found in actual Hagen–Poiseuille pipe flow. A direct instanton computation through action minimisation in plane Poiseuille flow can be extremely complex, and only instantons linking the laminar baseflow to non linear Tolmien–Schlichting type waves have been performed so far [Wan, 2013, Wan et al., 2015].

### v2

@@ 1.1 Bistability in physics and its description by statistical physics
Many physical systems, after they underwent a subcritical bifurcation or a first order transition (or a similar transition in more complex systems), display two or more metastable states. This situation means that the physical system can switch from one metastable state to another if it undergoes some statistically steady forcing, usually by crossing a boundary of the unforced dynamics separating the basin of attraction of one state from another. Conversely, this physical system will relax toward either states if it is prepared with an initial condition on the corresponding side of the separatrix, when no forcing noise is applied. Traditional bistable systems are found in kinetic chemistry Onsager, (1938); Hänggi et al., (1990); Van Erp et al., (2003); Lopes and Lelièvre, (2019), where the two states are the reactants and the products or different configurations or conformations Lopes and Lelièvre, (2019) of a complex molecule, while the force driving transitions is the thermal noise. Much insight on bistability has been gained from formal noise driven systems Metzner et al., (2006); Cérou et al., (2011); Bouchet and Touchette, (2012); Rolland and Simonnet, (2015); Rolland et al., (2016); Lucente et al., (2022). More recently, a lot of effort has also been directed toward the study of bistable climate systems Lucarini and Bódai, (2017, 2019); Baars et al., (2021); Herbert et al., (2020) and bistable turbulent flows Podvin and Sergent, (2017) of aerodynamical Grandemange et al., (2013); Kim and Durbin, (1988) or geophysical interest Berhanu et al., (2007); Bouchet et al., (2019). In these systems, the metastable states are usually large scale coherent circulations, like jets, or vortices, while the force driving transition is intrinsic to the system and related to the fluctuating small scales of turbulence.
There are several points of view to describe multistability arising from a deterministic system to which a random forcing is added. In this text we describe multistability in flows where turbulence can be sustained even in the absence of forcing: there are some differences with the case where turbulence is a consequence of said forcing Bouchet et al., (2019). Strictly speaking, the flow to which noise is added becomes a random dynamical system, to which the notions of stable and unstable attractors of autonomous dynamical systems cannot be applied Arnold et al., (1995). What is instead attracting the dynamics is a more complex object that represents the probable random fluctuations around either multistable states, with excursions from one multistable state to another. The application of such descriptions to small degrees of freedom systems has already been performed Chekroun et al., (2011). However, using such a point of view to describe multistable turbulent flows, with many degrees of freedoms is extremely complex and outside the scope of this text. Instead, we choose to follow an approach inspired by statistical physics, which has become popular as it is much more tractable. This requires the variance of the noise added to the deterministic part of the dynamics to be small enough. In that case, the properties of the unforced original system strongly determine those of the stochastically forced system. The first of these properties are the attractors of the unforced systems (or what is closest to attractors). Under the forcing of vanishing variance, the flow spends most of its time fluctuating near these attractors: the bistable states can then be defined as the narrow regions of phase space within one or two standard deviation of the attractors.
@@
From that point of view of statistical physics, a central object of study for metastable systems is the succession of states visited by the system when it transits from one bistable state to another. In kinetic chemistry, this part of the dynamics is termed a reactive trajectory Hänggi et al., (1990) or a transition path Van Erp et al., (2003). In kinetic chemistry, and more generally, in a wide class of stochastic systems, as the thermal noise is decreased and the transition becomes rare, the transition paths tend to concentrate around a specific path termed the instanton. Again, the phase space of the unforced deterministic system strongly influences the path followed by the instanton. The instanton generally consists in a fluctuation path under noise from the first attractor of the unforced dynamics toward a specific transition state which is a saddle of the deterministic part of the dynamics, followed by a deterministic relaxation path from that saddle to the second attractor Touchette, (2009) (Fig. 17 (a)). This path is important because it highlights the physical mechanisms of the transition, while the properties of the transition state control the parametric dependence of the transition rate Hänggi et al., (1990); Touchette, (2009); Bouchet and Reygner, (2016). While the instanton always visits the transition state, complex trajectories can be found in non-gradient systems Wan et al., (2015). The rarity and the concentration of trajectories is quantitatively studied in the case where noise variance is a constant times $\epsilon$, a vanishing parameter. This rarity is quantified by the the transition rate, which is the inverse of the mean first passage time before a transition occurs $T$. The mean first passage time goes to infinity exponentially with $\epsilon$, as
@@
We introduced the rate function $\mathcal{I}$. This quantity is finite (the limit is defined), not everywhere zero and is independent on $\epsilon$. This rate function is directly controlled by the properties of the escaped multistable state and the saddle. In kinetic chemistry, for relatively low temperatures, this relation was identified as Arrhenius law. It was then formalised in the context of stochastic gradient systems, systems where the deterministic part of the dynamics is the gradient of some potential $V$, as Eyring–Kramers theory Hänggi et al., (1990). In that case $\mathcal{I}$ is the different of potential between the first metastable state and the transition state: the saddle. Such a formula can also be obtained in non gradient systems Bouchet and Reygner, (2016). In that case, these exists a quasipotential that controls part of the dynamics, and $\mathcal{I}$ is the difference of quasipotential between the first metastable state and the saddle. In general, the asymptotic study of exponential dependences in probabilities or rate of probabilities (Eq. (1)) is termed Large Deviations, and encompasses the case of rare reactive trajectories. In the case of rare paths, this Large Deviation principle is named after Freidlin and Wentzell Freidlin and Wentzell, (1998). The difficulty of computing the transition path directly from the dynamics, either because of its complexity or because direct simulation times are too long for molecular dynamics, has led to the development and application of rare event simulation methods for kinetic chemistry Van Erp et al., (2003); Lopes and Lelièvre, (2019). Another approach for these computations consists in exploiting the fact that the instanton minimises the so-called Onsager–Machlup action Onsager and Machlup, (1953) in optimisation procedures Grafke and Vanden-Eijnden, (2019); Wan, (2013); Wan et al., (2015); Wan and Yu, (2017); Börner et al., (2023). Note that the noise covariance matrix is involved in the action, so that the type of forcing can influence the fluctuation path and the transition rate. The transition state, however, is a property of the deterministic part of the dynamics and thus has a degree of universality.
To conclude this section, we note that the addition of a noise makes bistability possible in systems where several stable fixed points can exist. In the case of low noise variance, statistical physics can greatly help the study of the problem. Among other things, it tells us that the properties of the noise induced bistability is strongly influenced by the phase space of the unforced system: the stable fixed points become the multistable states and the saddle between them will play a key role in the noise induced transitions. This means that it is very relevant to mention them when discussing reactive trajectories. It can also be very enlightening to present those path on top of the phase space of the unforced system.
1.2 Towards the use of rare events simulation methods in transitional wall turbulence
@@
The success of rare event simulation methods in chemistry has encouraged their use in fluid dynamics. For that purpose, they have been applied to spatially extended systems of gradual complexity, starting from one dimensional stochastically forced systems Rolland, (2018); Baars et al., (2021), moving to two dimensional stochastically forced turbulence Bouchet et al., (2019) and recently reaching three dimensional deterministic transitional turbulence Rolland, (2022); Gomé et al., (2022). The progress has been milestoned in this manner because of the size of the leap between stochastic differential equation and fully developed turbulent flows. This means a drastic increase of the number of degrees of freedom (DoF) involved, a complete change of the force driving the metastability, and sometimes a strong increase in complexity of the transition state between the two metastable states. While the extended systems studied were firstly chosen for their physical interest, the order of studies was also motivated by the need to tackle these changes separately from one another at first.
@@
This was the case of the most recent works, concerning the collapse of turbulence in plane Couette flow Rolland, (2022); Gomé et al., (2022). Plane Couette flow (PCF) is the flow between two parallel moving walls (Fig. 1, left). In PCF, like in other wall flows such as plane Poiseuille flow Wan, (2013); Wan et al., (2015); Wan and Yu, (2017), Couette–Poiseuille flow Liu et al., (2021), Hagen–Poiseuille pipe flow Willis and Kerswell, (2009) or boundary layers Spangler and Wells Jr, (1968); Rigas et al., (2021), wall turbulence can coexist at Reynolds numbers (or positions in the case of boundary layers) for which the laminar baseflow is linearly stable (see Romanov, (1973) for the case of PCF). As a consequence, transitional turbulence can collapse under its own fluctuations, while the flow can go from laminar to turbulent under a forcing Spangler and Wells Jr, (1968); Rigas et al., (2021) or if it is given an initial condition “close enough” to turbulence Schmiegel and Eckhardt, (1997). The study of the collapse of turbulence using rare event methods gave the occasion to investigate the physical mechanism of laminar hole opening as well as the technical solutions to study rare multistability events in a three dimensional flow not stochastically forced Rolland, (2022); Gomé et al., (2022). Conversely, the study of the forcing of laminar flow toward turbulence, under a noise or another, would give the occasion to investigate the crossing of a very complex separatrix  [^1] We use the term separatrix for lack of a better word. strictly speaking, an unforced wall flow starting from any initial condition eventually reaches the laminar state: what changes is the relaxation time before it does so. This duration dramatically increases as an hypersurface is crossed Schmiegel and Eckhardt, (1997)..
@@
As such, this has brought the question of the separatrix between laminar and turbulent flow at the forefront of the study of transition. The first descriptions by Waleffe and collaborators of the mechanisms maintaining transitional wall turbulence was the self sustaining process (SSP) of wall turbulence Hamilton et al., (1995); Waleffe, (1997). This process represented wall turbulence as a stable limit cycle, which was therefore separated from the laminar baseflow in phase space by an actual separatrix with a saddle on it. The SSP describes how streamwise velocity streaks and streamwise vortices regenerate one another. It can take place provided the Reynolds number is large enough and the domain size is larger than the so called Minimal Flow Unit (MFU) Jiménez and Moin, (1991). Refinements of that picture indicated that this process was not periodic nor an attractor (thus leaving the room for collapse of turbulence) but nevertheless confirmed the idea that there was a separatrix, on which saddles could be found. The structure of the separatrix, which is a key matter for the forced path from laminar to turbulent flow, was often studied using methods of dichotomy or bisections Nusse and Yorke, (1989). This has led to the study of edge states, saddles on that separatrix, in wall flows Toh and Itano, (2003); Schneider et al., (2007, 2008, 2010); Willis and Kerswell, (2009) as well as in climate models Lucarini and Bódai, (2017). Note that additional unstable limit cycles and fixed points can be found on the turbulent side of the separatrix Gibson et al., (2008). In wall flows, these unstable fixed points and limit cycles are often reminiscent of the turbulent flow, in that they display velocity streaks and streamwise vortices. However, they display less small scale fluctuations and are often very spatially regular. This regularity can be characterised using the symmetry groups of the flow Gibson et al., (2008); Schneider et al., (2010). Note finally that, as the separatrix and the unstable states on it are a property of the flow in the configuration of interest and not of initial conditions or forcing type, they also represent an universal feature of transition to turbulence.
@@
In wall flows, the complexity of this boundary is such that the saddle on it is not the point of the separatrix which is has the smallest amplitude of departure to the laminar baseflow. Experimental and numerical studies of initial conditions, generated using various specific rules, of minimal energy that evolve into turbulence have highlighted that. This leaves the question of whether the most probable forced path would visit this saddle at very small but finite forcing noise variance. Alternate investigations have thus been aimed at identifying a path forced and/or triggered by a finite amplitude initial condition that would drive the flow from laminar to turbulent. Linear studies of optimal initial conditions of infinitesimal amplitude leading to maximal amplification of energy identified lift-up: the extraction of energy from the laminar base flow from streamwise vortices. The lift-up mechanism was thus identified in transient growth studies that highlight the fact that energy growth can happen while all linear modes are stable if the linearised Navier–Stokes equations lead to a non normal operator Farrell and Ioannou, 1993a ; Schmid and Henningson, (1994); Trefethen et al., (1993). Such an optimal (in shape) initial condition whose amplitude places it on the laminar side of the boundary would remain on the laminar side of said boundary. However, for initial conditions starting on the turbulent side of the boundary, lift-up is an important part of the self-sustaining process Waleffe, (1997). Refinement of optimisation of initial condition energy amplification technics can be used to compute the optimal finite amplitude initial condition leading to maximal non-linear energy growth Cherubini et al., (2011); Rabin et al., (2012); Monokrousos et al., (2011). This initial condition is termed a minimal seed. Such computations are still performed on the laminar side of the separatrix, however a slight modification of the seed, in the right direction, can generate an initial condition which has very little energy but can nevertheless travel to turbulence extremely efficiently. Note that by discretising this procedure in phase space, the instanton linking different stable fixed points can be computed Lecoanet and Kerswell, (2018). There is in practice no explicit stochastic forcing in that case, but the resulting trajectory still undergoes a fluctuation toward the relevant saddle, followed by a relaxation on which forcing is absent.
@@
The question of whether a random forcing of a wall flow can generate a minimal seed, a flow organisation whose shape and amplitude is precisely defined, remains open. Indeed, the study of forced wall flows gives little weight to effects like the Orr mechanism, involved in the growth of minimal seeds. In parallel with studies of optimal initial conditions leading to transient growth, the study of the response of the linearised Navier–Stokes equations to a forcing with a wide range of spatial spectra, which are white in time, identified a reacting flow structure Farrell and Ioannou, 1993b . The weakly non linear behaviour of flows under such stochastic forcing can also be performed Ducimetière et al., (2022). The inclusion of the strong non-linearities may very well strongly change this picture because they introduce the aforementioned complex separatrix. The experimental study of the response of boundary layers to harmonic forcing of different spectrum has a long history and shows a strong dependence on the forcing frequency and amplitude of the point of appearance of wall turbulence Spangler and Wells Jr, (1968). Numerical equivalent of such studies of the response to an harmonic forcing have only been recently undertaken Rigas et al., (2021). Meanwhile, to refocus on our subject of interest, studying the response to a stochastic forcing of vanishing variance requires a rare event simulation method. This has been performed in the case of a model of Hagen–Poiseuille pipe flow Rolland, (2018). In that case, we have termed the forcing of the flow model from laminar to turbulent build up. However the separatrix in that model is nowhere near as complex as the one found in actual Hagen–Poiseuille pipe flow. A direct instanton computation through action minimisation in plane Poiseuille flow can be extremely complex, and only instantons linking the laminar baseflow to non linear Tolmien–Schlichting type waves have been performed so far Wan, (2013); Wan et al., (2015).
1.3 Outline of the article